# Structurally Flexible Neural Networks: Evolving the Building Blocks for General Agents

### v1

@@ 5. Results
Figure 6. Lifetime Progress in Training Environments: Columns display the results under different conditions in each of the three environments. The first three columns of subplots show the results of the agent of the best of the five evolution runs of the standard SFNN approach, described in Section 3. The three remaining columns show the results of an agent trained on a single, fixed adjacency matrix throughout its evolution (Fixed_SFNN). Each bar corresponds to the score in one episode of the lifetime of an agent. Blue: Adjacency matrices are initialized randomly at the beginning of the agent’s lifetime. Green: The same adjacency is used in each lifetime. This is the same adjacency matrix used in the evolution of Fixed_SFNN. Yellow: Adjacency matrices are initialized randomly and random permutations of inputs and outputs are chosen at the beginning of the agent’s lifetime. The SFNN optimized under standard conditions shows similar lifetime performances across all three settings. The SFFN that was only exposed to a single adjacency matrix during its evolution obtains good scores when that specific matrix is used (green), but exhibits major performance decreases in the other two settings. The shown scores are the averaged scores across 100 lifetimes.
In Figure 6, average lifetime scores of an evolved SFNN over 100 lifetimes can be found. Scores of the other variations from Figure 5 are omitted, as no good solutions were found by evolution for any of them. In the first three columns, the SFNN is evaluated in the same setting as during the training, in a setting with the same adjacency matrix used in all lifetimes, and in a setting where a random permutation of the inputs and outputs are chosen at the beginning of each lifetime. In all cases, the scores of the episodes later in the agent’s life tend to be better, indicating that the networks in a short amount of time are able to organize into functional networks.
@@ 6. Discussion and Future Work
Of the four settings tested in Figure 5, only the SFNN with different neuron and synapse types and sparse random connectivity was able to improve on all tasks. The SymLA approach with its relatively few plastic parameters might lack the capacity to form networks that are different enough to adapt to multiple tasks. The SFNN_single, on the other hand, is potentially under-parameterized in terms of the evolved parameters. SFNN_fully was the worst scoring setting of them all, hinting that having different neuron and synapse types alone is not enough to overcome the symmetry dilemma: a fully connected structure has a high risk of ending up with all units having the same activity, due to each unit getting the same input.
@@
Another option could be to use Set Transformers that are specifically designed to be permutation invariant and would not rely on being exposed to all permutations during the training phase (Lee et al., 2019). To make this type of model invariant to the length of the observation vectors from different environments, the set would be the observation elements for a given environment, instead of sequences of multiple time steps. However, the standard architecture of Set Transformers does include fully connected feedforward layers in the network. This limits the possibility of varying the size of the networks after the evolution phase. While not part of the experiments shown in this paper, true structurally flexible neural networks should also be able to have varying hidden layer sizes across different lifetimes, enabling a trade-off between redundancy and stability on one hand, and computational efficiency on the other.
@@
The experiments in this paper used environments with small input and output sizes. Challenges related to the symmetry dilemma are magnified when input and output sizes are increased. More specifically, more input elements potentially make it harder for the input synapses with shared parameters to distinguish them and make properly individualized synaptic strengths for each of them. More experiments are needed to determine whether more types of neurons and synapses in the hidden reservoir are sufficient to mitigate the symmetry dilemma in high-dimensional environments.
Breaking the symmetry of the network by placing neurons in the same layer with a random adjacency matrix requires the training phase to incorporate different permutations of the adjacency matrix and ordering of neurons so that the evolved parameters do not overfit to any specific configuration. Further, looking at the second column Figure 6, the SFNN shows better average performances in two out of three environments, when a single adjacency matrix was used in all lifetimes. This points towards some adjacency matrices being easier to adapt to than others. With these points in mind, instead of relying on random connectivity future studies could potentially benefit from combining the SFNN approach with an algorithm for structural development (Najarro et al., 2023).
@@
Artificial neural networks that can be optimized across different environments to quickly adapt their synaptic weights to form a functional network have the prerequisites to develop into a general learner. With this paper, we contribute to this direction by showing promising results with the introduced SFNN approach. With a relatively small number of evolved parameters, an SFNN improves during its lifetime on three different control tasks. We believe that scaling up this and related approaches to many more tasks has the potential to yield new exciting opportunities in the field of artificial agents.

### v2

@@ 5. Results
Figure 5. Adapting Weight Matrices in SFNNs. An example of the same initial weight matrix used in the three environments resulting in different weight matrices after eight episodes in a different task. The sum of the four elements that make up the synapse is depicted. Values are clipped to have a maximum magnitude of 10 for readability. The only difference in the initial matrices between the environments is how many neurons are counted as input/output neurons, to fit the respective environment. Especially the solution for the CartPole environment stands out from the two others with the magnitudes of weights coming from the input neurons being much larger than the other weights.
Figure 6. Lifetime Progress in Training Environments. Columns display the results under different conditions in each of the three environments. The shown scores are the averaged scores across 100 lifetimes. The first three columns of subplots show the results of the agent of the best of the five evolution runs of the standard SFNN approach, described in Section 3. The three remaining columns show the results of an agent trained on a single, fixed adjacency matrix throughout its evolution (Fixed_SFNN). Each bar corresponds to the score in one episode of the lifetime of an agent. Blue: Adjacency matrices are initialized randomly at the beginning of the agent’s lifetime. Green: The same adjacency is used in each lifetime. This is the same adjacency matrix used in the evolution of Fixed_SFNN. Yellow: Adjacency matrices are initialized randomly and random permutations of inputs and outputs are chosen at the beginning of the agent’s lifetime. The SFNN optimized under standard conditions shows similar lifetime performances across all three settings. The SFFN that was only exposed to a single adjacency matrix during its evolution obtains good scores when that specific matrix is used (green), but exhibits major performance decreases in the other two settings. In sum, the standard SFNN approach can improve quickly across settings, while the SFNN trained on a single adjacency matrix cannot.
Figure 5 shows an example of how one initial weight matrix can diverge into different ones when the network is tasked with solving one environment or the other. This shows that the dynamics of the single set of evolved neurons and synapses can self-organize into several configurations that are adaptive to the given environment.
Figure 6 shows average lifetime scores of an evolved SFNN over 100 lifetimes. Scores of the other variations from Figure 4 are omitted, as no good solutions were found by evolution for any of them. In the first three columns, the SFNN is evaluated in the same setting as during the training, in a setting with the same adjacency matrix used in all lifetimes, and in a setting where a random permutation of the inputs and outputs are chosen at the beginning of each lifetime. In all cases, the scores of the episodes later in the agent’s life tend to be better, indicating that the networks in a short amount of time are able to organize into functional networks.
@@ 6. Discussion and Future Work
Of the four settings tested (Figure 4), only the SFNN with different neuron and synapse types and sparse random connectivity was able to improve on all tasks. The SymLA approach with its relatively few plastic parameters might lack the capacity to form networks that are different enough to adapt to multiple tasks. The SFNN_single, on the other hand, is potentially under-parameterized in terms of the evolved parameters. SFNN_fully was the worst scoring setting of them all, hinting that having different neuron and synapse types alone is not enough to overcome the symmetry dilemma: a fully connected structure has a high risk of ending up with all units having the same activity, due to each unit getting the same input.
@@
Another option for more flexible neural networks could be the use of Set Transformers that are specifically designed to be permutation invariant and would not rely on being exposed to all permutations during the training phase (Lee et al., 2019). To make this type of model invariant to the length of the observation vectors from different environments, the set would be the observation elements for a given environment, instead of sequences of multiple time steps. This is similar to having parameter-sharing input neurons and synapses as with SFNNs. The attention matrix of the Transformer can be interpreted as plastic parameters (Schlag et al., 2021). However, whereas the size of the attention matrix of the Transformer is tied to the number of input elements, SFNNs allow the number of plastic parameters to be independent of the number of input elements. While not part of the experiments shown in this paper, true structurally flexible neural networks should also be able to have varying hidden layer sizes across different lifetimes, enabling a trade-off between redundancy and stability on one hand, and computational efficiency on the other.
@@
The experiments in this paper used environments with small input and output sizes. Challenges related to the symmetry dilemma are magnified when input and output sizes are increased. More specifically, more input elements potentially make it harder for the input synapses with shared parameters to distinguish them and make properly individualized synaptic strengths for each of them. More experiments are needed to determine whether more types of neurons and synapses in the hidden reservoir are sufficient to mitigate the symmetry dilemma in high-dimensional environments. Similarly, future experiments should examine the SFNN approach in environments with continuous action spaces. Since choosing discrete actions, as in the environments used here, forcefully breaks symmetry among output neurons’ signals, continuous action spaces might prove more challenging.
Breaking the symmetry of the network by placing neurons in the same layer with a random adjacency matrix requires the training phase to incorporate different permutations of the adjacency matrix and ordering of neurons so that the evolved parameters do not overfit to any specific configuration. Further, looking at the second column in Figure  6, the SFNN shows better average performances in two out of three environments, when a single adjacency matrix was used in all lifetimes. This points towards some adjacency matrices being easier to adapt to than others. With these points in mind, instead of relying on random connectivity future studies could potentially benefit from combining the SFNN approach with an algorithm for structural development (Najarro et al., 2023).
@@
Artificial neural networks, which can be optimized across different environments to quickly adapt their synaptic weights to form a functional network, have the prerequisites to develop into a general learner. Through the promising results demonstrated with the introduced SFNN approach in this paper, we hope to contribute to this direction. With a relatively small number of evolved parameters, an SFNN improves during its lifetime on three different control tasks. We believe that scaling up this and related approaches to many more tasks has the potential to yield new exciting opportunities in the field of artificial agents.